# ReeM: Ensemble Building Thermodynamics Model for Efficient HVAC Control via Hierarchical Reinforcement Learning

### v1

@@ 1 Introduction
Recently, the control optimization of HVAC systems has evolved from traditional methods (e.g., bang-bang control, feedback-based control) to Model Predictive Control (MPC), a constrained optimal control strategy that determines the optimal control inputs by minimizing a predefined objective function Yao and Shekhar (2021). For instance, an electrical price-aware MPC scheme for floor heating systems was developed for a residential apartment in Denmark Hu et al. (2019). As the name suggests, the MPC scheme requires appropriate modeling of systems and components to effectively control the dynamic processes within the building and HVAC system. The model serves as the cornerstone of the MPC system, enabling the prediction of future thermodynamics—typically reflected by indoor temperature [^1] We emphasize that predicting thermodynamics is analogous to predicting indoor temperature, a practice commonly found in the literature.—of the building environment. A schematic diagram of MPC in building HVAC systems is presented in Fig 1. Various modeling schemes have been developed for the associated environments, including both physics-based and data-driven approaches Yao and Shekhar (2021). With a more accurate thermodynamic model, the effectiveness of control actions derived from optimization can be significantly enhanced.
Figure 1: a) The process of HVAC control optimization; b) To search the optimal control at time $t$, the model conducts what-if predictions.
@@ 2 Background and Related Work
$f(\cdot,\cdot,\cdot)=\underbrace{Ax_{t-}+B_{1}u_{(t-1)-}+Ed_{t-}}_{\text{observed part}}+\underbrace{B_{2}u_{t}}_{\text{control variable}}$ (2)
@@ 3 Problem Statement
To model a given environment $e$, existing strategies typically develop a specific function $f_{e}$ for that environment, as illustrated in Fig 2(a). In this paper, we revisit this problem from a model ensemble perspective, as shown in Fig 2(b). Specifically, given the features $\mathbf{x_{t}}$, $\mathbf{u_{t-1}}$, and $\mathbf{d_{t}}$ collected at each time step $t$, along with a set of $N$ customized models $\{f_{i}\}_{i=1}^{N}$ developed for diverse building environments, our goal is to automatically construct a model ensemble $f^{ens}_{t}$ to achieve the best prediction performance by determining the optimal weight vector $\mathbf{w}=[w_{1},...,w_{N}]$ for the base models. The ensemble model is defined as: $f_{t}^{ens}=\sum_{i=1}^{N}w_{i}f_{i}$ and the predicted state at time $t+1$ is given by $\hat{x}_{t+1}=f_{t}^{ens}(\cdot,\cdot,\cdot)$ when input a control $u_{t}$ at time $t$. The problem can be formulated as: $\mathbf{w}^{*}=\arg\min_{\mathbf{w}}\mathcal{L}(f_{t}^{ens})$, where $\mathcal{L}$ denotes the prediction error.
Figure 2: Paradigms of preparing the model for MPC application. (a) modeling a customized model. (b) ensemble model via the developed models.
@@ 4.2 Encoder for State Representation
Before the RL design, we first introduce a state encoder module that focuses on extracting useful data representations as the state for the following agents. The encoder architecture is illustrated in the left part of Fig 3 with input information from both building data and base models. For building data, we adopt temporal convolutional networks (TCN) Bai et al. (2018) for representation learning. The networks consist of a stack of three TCN blocks with a kernel size of 4 and dilation rates of 1, 2, and 4. This architecture is designed to ensure that the network possesses a sufficiently long receptive field and each layer captures temporal patterns at varying granularities. Additionally, to better capture the interrelationships, we introduce a cross-attention module for effective feature fusion. Specifically, we set: $\mathbf{q}=\mathbf{H_{q}W_{q}},\mathbf{k}=\mathbf{H_{k}W_{k}},\mathbf{v}=\mathbf{H_{v}W_{v}}$. Here, $\mathbf{H_{q}}=\tau_{x}(\mathbf{x_{t}})$ and $\mathbf{H_{k}}=\mathbf{H_{v}}=\tau_{u}(\mathbf{u_{t-1}})\oplus\tau_{d}(\mathbf{d_{t}})$, where $\tau$ represent the TCN networks. $\mathbf{W_{q}}\in\mathbb{R}^{d_{h}\times d_{h}}$, $\mathbf{W_{k}}$ and $\mathbf{W_{v}}\in\mathbb{R}^{(2d_{h})\times d_{h}}$ are learnable projection matrices for attention. Then, the embedding is computed as: $e_{t}^{d}=\text{Softmax}(\mathbf{qk}^{T}/\sqrt{d_{h}})\mathbf{v}$, where $d_{h}$ is the hidden dimension. Furthermore, we evaluate the prediction performance of base models on historical data as supplementary information for model selection. Specifically, the mean squared error of each model at the last timestamp $t-1$ is recorded and embedded as: $e_{t}^{m}=\text{Embed}(\{\mathcal{L}_{t-1}(f_{i})\}_{i=1}^{N})$. Thus, the state representation is defined as:
@@ 4.3 High-level Agent
Action. We define the high-level action $a_{t}^{h}$ as a vector of binary values $[b_{1},...,b_{N}]$, where $b_{i}=1$ indicates that the model $f_{i}$ is selected. Conversely, if $b_{i}=0$, $f_{i}$ will not be considered in the model ensemble.
Reward. The reward should indicate whether the model ensemble performs effectively, thereby guiding the agent’s learning process. Consequently, an intuitive reward is the forecasting error of the ensemble model $f_{t}^{ens}$ at time $t$, defined as $r_{t}^{loss}=-\mathcal{L}_{t}(f_{t}^{ens})$. However, this reward design alone may lead the high-level agent to select an excessive amount of irrelevant models, which increases the training burden on the low-level agent. Besides, the model ensemble $f_{t}^{ens}$ derived from an extensive combination of models is unlikely to accurately reflect the real building environment due to the complex factors and features involved Picard et al. (2017). To encourage the high-level agent to select a reasonably smaller amount of models with simpler structures, we introduce two internal rewards $r_{t}^{mod}=-\sum_{i=1}^{N}b_{i}$ and $r_{t}^{var}=-||\sum_{i=1}^{N}b_{i}f_{i}||$, which penalize the number of base models and variables selected. Here |||||$\cdot$||| | ⋅ | | measures the amount of variables in a model. Then, the high-level reward can be represented as:
@@ 4.4 Low-level Agent
Action. We define the low-level action $a_{t}^{l}$ as a vector of weights $[w_{1},...,w_{N}]$, where $w_{i}=0$ if $b_{i}=0$ and $w_{i}\in[0,1]$ if $b_{i}=1$. Since the indoor temperature typically remains within a relatively consistent range across buildings, and given that the model library contains sufficiently diverse base models, we enforce the constraint $\sum_{i=1}^{N}w_{i}=1$ to narrow the action space for efficient weight search without significantly compromising prediction accuracy.
@@ 5.1 Evaluation Methodology
Figure 4: The performance of the base models applied to the test rooms, measured by the Mean Absolute Error (MAE) metric.
Base Models Preparation. As shown in Fig 2(b), the ensemble approaches rely on a model library that comprises a set of base models. For the preparation of these base models, we follow the traditional paradigm depicted in Fig 2(a) We employ two existing data-driven modeling methodologies, resulting in two customized models for each room: (i) Symbolic Regression (SR): This method aims to identify a mathematical expression to represent the relationships between variables and is commonly used in this context, such as for one-day-ahead prediction Leprince et al. (2021) and 15-minute-ahead prediction Zhao et al. (2023). (ii) Multiple Linear Regression (MLR): A more classical modeling approach in thermodynamics, which we re-implement  Moretti et al. (2021) where the variables involved are the installed sensors and day type. Fig 4 presents the results of directly applying these base models. Notably, no single model consistently performs well across different rooms over the two-day evaluation.
@@
Baselines. We compare ReeM with the following baselines: (1) Heuristics method. We first compare ReeM with a classical heuristic method commonly used in dynamic model ensemble problems in data streaming scenarios Cerqueira et al. (2017). We implement a simple scheme that leverages the best $N$ models at previous time step $t-1$ as the ensemble, using average weighting. In our experiments, $N=3$. (2) RLMC Fu et al. (2022). This is an RL-based scheme utilizing DDPG, an actor-critic algorithm for dynamic model combination in time-series forecasting. (3) DMS Feng and Zhang (2019). This is another Q-learning-based RL scheme for dynamic model selection. (4) BOA Deng et al. (2024). We also compare ReeM with an ensemble method in the building energy domain. Unlike the aforementioned RL-based ensemble methods, BOA generates a static ensemble formulated as a Bayesian optimization problem and uses Gaussian processes (GP) to search the optimal combination.
@@
Figure 5: The comparison of five ensemble methods in the 13 tested rooms. Values interms of the MAE metric.
@@ 5.3 The Ablation Study (Q2)
Figure 6: Heating experimental results: (a) outdoor temperature, (b) indoor temperature and HVAC temperature setpoint, (c) HVAC unit power consumption, (d) HVAC unit energy consumption. SR-based MPC and proposed ReeM-based HVAC management framework were conducted on February 8 and 9.
Figure 7: The key parameters setting.
@@ 5.4 The Study of Key Hyperparameters (Q3)
[…]
@@ 6 On-site Case Study
As shown in Fig. 6(b), both SR and ReeM effectively control and maintain the room temperature within the required range during the target period. Additionally, Fig. 6(c) and (d) indicate that the average HVAC power consumption of ReeM is slightly lower than that of the customized SR under an automation process. These experimental results confirm that the ReeM approach can provide an accurate and efficient thermodynamic model for HVAC operation optimization. Moreover, ReeM enables real-time online updates of the thermodynamic model without requiring re-training, a breakthrough for dynamic systems where adaptability is critical, yet SR model updates typically involve computationally expensive retraining, which struggles to keep up with rapid environmental changes.

### v2

# Towards Machine Learning-based Model Predictive Control for HVAC Control in Multi-Context Buildings at Scale via Ensemble Learning

@@ 1 Introduction
Recently, the control optimization of HVAC systems has evolved from traditional methods (e.g., bang-bang control, feedback-based control) to Model Predictive Control (MPC), a constrained optimal control strategy that determines the optimal control inputs by minimizing a predefined objective function Yao and Shekhar (2021). For instance, an electrical price-aware MPC scheme for floor heating systems was developed for a residential apartment in Denmark Hu et al. (2019). As the name suggests, the MPC scheme requires appropriate modeling of systems and components to effectively control the dynamic processes within the building and HVAC system. The model serves as the cornerstone of the MPC system, enabling the prediction of future thermodynamics—typically reflected by indoor temperature [^1] We emphasize that predicting thermodynamics is analogous to predicting indoor temperature, a practice commonly found in the literature.—of the building environment. A schematic diagram of MPC in building HVAC systems is presented in figure. Various modeling schemes have been developed for the associated environments, including both physics-based and data-driven approaches Yao and Shekhar (2021). With a more accurate thermodynamic model, the effectiveness of control actions derived from optimization can be significantly enhanced.
@@ 2 Background and Related Work
[…]
@@ 3 Problem Statement
To model a given environment $e$, existing strategies typically develop a specific function $f_{e}$ for that environment, as illustrated in Fig LABEL:fig:_story(a). In this paper, we revisit this problem from a model ensemble perspective, as shown in Fig LABEL:fig:_story(b). Specifically, given the features $\mathbf{x_{t}}$, $\mathbf{u_{t-1}}$, and $\mathbf{d_{t}}$ collected at each time step $t$, along with a set of $N$ customized models $\{f_{i}\}_{i=1}^{N}$ developed for diverse building environments, our goal is to automatically construct a model ensemble $f^{ens}_{t}$ to achieve the best prediction performance by determining the optimal weight vector $\mathbf{w}=[w_{1},...,w_{N}]$ for the base models. The ensemble model is defined as: $f_{t}^{ens}=\sum_{i=1}^{N}w_{i}f_{i}$ and the predicted state at time $t+1$ is given by $\hat{x}_{t+1}=f_{t}^{ens}(\cdot,\cdot,\cdot)$ when input a control $u_{t}$ at time $t$. The problem can be formulated as: $\mathbf{w}^{*}=\arg\min_{\mathbf{w}}\mathcal{L}(f_{t}^{ens})$, where $\mathcal{L}$ denotes the prediction error.
@@ 4.2 Encoder for State Representation
Before the RL design, we first introduce a state encoder module that focuses on extracting useful data representations as the state for the following agents. The encoder architecture is illustrated in the left part of figure with input information from both building data and base models. For building data, we adopt temporal convolutional networks (TCN) Bai et al. (2018) for representation learning. The networks consist of a stack of three TCN blocks with a kernel size of 4 and dilation rates of 1, 2, and 4. This architecture is designed to ensure that the network possesses a sufficiently long receptive field and each layer captures temporal patterns at varying granularities. Additionally, to better capture the interrelationships, we introduce a cross-attention module for effective feature fusion. Specifically, we set: $\mathbf{q}=\mathbf{H_{q}W_{q}},\mathbf{k}=\mathbf{H_{k}W_{k}},\mathbf{v}=\mathbf{H_{v}W_{v}}$. Here, $\mathbf{H_{q}}=\tau_{x}(\mathbf{x_{t}})$ and $\mathbf{H_{k}}=\mathbf{H_{v}}=\tau_{u}(\mathbf{u_{t-1}})\oplus\tau_{d}(\mathbf{d_{t}})$, where $\tau$ represent the TCN networks. $\mathbf{W_{q}}\in\mathbb{R}^{d_{h}\times d_{h}}$, $\mathbf{W_{k}}$ and $\mathbf{W_{v}}\in\mathbb{R}^{(2d_{h})\times d_{h}}$ are learnable projection matrices for attention. Then, the embedding is computed as: $e_{t}^{d}=\text{Softmax}(\mathbf{qk}^{T}/\sqrt{d_{h}})\mathbf{v}$, where $d_{h}$ is the hidden dimension. Furthermore, we evaluate the prediction performance of base models on historical data as supplementary information for model selection. Specifically, the mean squared error of each model at the last timestamp t1t-1 is recorded and embedded as: $e_{t}^{m}=\text{Embed}(\{\mathcal{L}_{t-1}(f_{i})\}_{i=1}^{N})$. Thus, the state representation is defined as:
@@ 4.3 High-level Agent
Action. We define the high-level action $a_{t}^{h}$ as a vector of binary values $b_{1},...,b_{N}]$, where $b_{i}=1$ indicates that the model $f_{i}$ is selected. Conversely, if $b_{i}=0$, $f_{i}$ will not be considered in the model ensemble.
[…]
@@ 4.4 Low-level Agent
Action. We define the low-level action $a_{t}^{l}$ as a vector of weights $w_{1},...,w_{N}]$, where $w_{i}=0$ if $b_{i}=0$ and $w_{i}\in[0,1]$ if $b_{i}=1$. Since the indoor temperature typically remains within a relatively consistent range across buildings, and given that the model library contains sufficiently diverse base models, we enforce the constraint $\sum_{i=1}^{N}w_{i}=1$ to narrow the action space for efficient weight search without significantly compromising prediction accuracy.
@@ 5.1 Evaluation Methodology
Base Models Preparation. As shown in figure, the ensemble approaches rely on a model library that comprises a set of base models. For the preparation of these base models, we follow the traditional paradigm depicted in figure. We employ two existing data-driven modeling methodologies, resulting in two customized models for each room: (i) Symbolic Regression (SR): This method aims to identify a mathematical expression to represent the relationships between variables and is commonly used in this context, such as for one-day-ahead prediction Leprince et al. (2021) and 15-minute-ahead prediction Zhao et al. (2023). (ii) Multiple Linear Regression (MLR): A more classical modeling approach in thermodynamics, which we re-implement  Moretti et al. (2021) where the variables involved are the installed sensors and day type. Figure presents the results of directly applying these base models. Notably, no single model consistently performs well across different rooms over the two-day evaluation.
@@
Baselines. We compare ReeM with the following baselines: (1) Heuristics method. We first compare ReeM with a classical heuristic method commonly used in dynamic model ensemble problems in data streaming scenarios Cerqueira et al. (2017). We implement a simple scheme that leverages the best $N$ models at previous time step t1t-1 as the ensemble, using average weighting. In our experiments, $N=3$. (2) RLMC Fu et al. (2022). This is an RL-based scheme utilizing DDPG, an actor-critic algorithm for dynamic model combination in time-series forecasting. (3) DMS Feng and Zhang (2019). This is another Q-learning-based RL scheme for dynamic model selection. (4) BOA Deng et al. (2024). We also compare ReeM with an ensemble method in the building energy domain. Unlike the aforementioned RL-based ensemble methods, BOA generates a static ensemble formulated as a Bayesian optimization problem and uses Gaussian processes (GP) to search the optimal combination.
@@ 5.4 The Study of Key Hyperparameters (Q3)
We explore the effects of the two key hyperparameters on the model’s performance. i) Effects of the Base Model Number. We adjusted the number of base models from 10, 20, and 30 up to 104 (the total number of models). The results presented in figure indicate that the proposed ReeM achieves optimal performance at 30 models, with performance declining from 40 models onward. We attribute this phenomenon to the relative insufficiency of training data for ReeM to effectively explore such a large action space. For comparison, we conducted the same study in RLMC, where performance decreased from 10 models. ii) Effects of the Type of base model. We vary the three types of the base models in $\{\text{SR},\text{MLR},\text{SR+ MLR}\}$ and compared their performance. The results are presented in figure. It was observed that the type of base model has minimal impact; specifically, the proposed ReeM does not heavily rely on any particular modeling method for the base model.
@@ 6 On-site Case Study
As shown in figure, both SR and ReeM effectively control and maintain the room temperature within the required range during the target period. Additionally, Fig. LABEL:fig:online-exp(c) and (d) indicate that the average HVAC power consumption of ReeM is slightly lower than that of the customized SR under an automation process. These experimental results confirm that the ReeM approach can provide an accurate and efficient thermodynamic model for HVAC operation optimization. Moreover, ReeM enables real-time online updates of the thermodynamic model without requiring re-training, a breakthrough for dynamic systems where adaptability is critical, yet SR model updates typically involve computationally expensive retraining, which struggles to keep up with rapid environmental changes.